The sum of two numbers is 62. The smaller number is 6 less than the larger number. What are the numbers?
Found 2 solutions by ewatrrr, joyesh29-03:
Answer by ewatrrr(24785)

(x-6) + x = 62
2x = 68
x = 34
checking...
28 + 34 = 62

Answer by joyesh29-03(1)

let one number be -x,and other be -y
then according to the question-x=y-6 equation 1
and x+y=62 equation 2
substituting equation 1 in equation 2
therefore y+y-6=62
2y=62+6
2y=68
y=34 equation 3
putting equation 3 in equation 1
x=34-6
x=28

therefore x=28 and y=34